OpenStudy (anonymous):

In autumn, leaves on a tree turn yellow and start to drop. When the first gust of wind blows, half of the leaves, plus an additional 10 leaves, fall. When the second gust of wind blows, half of the leaves that remain, plus an additional 10 leaves, fall. When the third goes of wind blows, half of the leaves that remain, plus an additional 10 leaves, fall. At this point, there are no more leaves on the tree. How many leaves are there on the tree before the first gust of wind came along?

OpenStudy (dumbcow):

shoot sorry my browser froze and had to refresh and lost everything i had written :{

OpenStudy (dumbcow):

anyway let x be total leaves set x equal to sum of fallen leaves in each gust \[x = (\frac{x}{2}+10)+(\frac{x-20}{4}+10)+(\frac{x-60}{8}+10)\]

OpenStudy (anonymous):

thank you @dumbcow :).

OpenStudy (dan815):